OpenStudy (anonymous):

10i/-5+5i still confused

OpenStudy (anonymous):

multiply top and bottom my conjugate

OpenStudy (anonymous):

me either, what happen with 10i/-5+5i ?

OpenStudy (amistre64):

\[\frac{10i}{-5} +5i??\]

OpenStudy (anonymous):

no the +5i is under the division sign too

OpenStudy (anonymous):

10i / (-5 + 5i) = 1 - i

OpenStudy (anonymous):

that is given i as an imaginary unit - what is the actual question? See if this helps given i as a variable = http://www.wolframalpha.com/input/?i=10i%2F%28-5%2B5i%29&a=i_Variable